A rectangle has an area of 36 square feet and a base of 9 feet . what is the perimeter of the rectangle?
Answer by macston(5194)

L=length=base; W=width; A=area=L*W; P=perimeter=2(L+W)
A=L*W
36 sq ft=9 ft*W Divide each side by 9 ft
4 ft=W The width of the rectangle is 4 feet.
P=2(L+W)=2(9 ft+4 ft)=2(13 ft)=26 ft
ANSWER: The perimeter of the rectangle is 26 feet.